OpenStudy (anonymous):

estimate the square root to the nearest integer and tenth squareroot 175

OpenStudy (anonymous):

does "tenth square root" mean "tenth root"?

OpenStudy (anonymous):

I dont have the squareroot sign so i just typed it..it is 175

OpenStudy (anonymous):

nearest integer and tenth

zepdrix (zepdrix):

Estimate? :o Hmmm is there some specific method for estimating? I can't seem to remember...

zepdrix (zepdrix):

175 is awful close to 169 which is a perfect square.

zepdrix (zepdrix):

\[\Large\bf\sf \sqrt{175}\quad=\quad \sqrt{169+6}\]Not really sure what we would do from there...

zepdrix (zepdrix):

Factoring 169 out of each,\[\Large\bf\sf \sqrt{169\left(1+\frac{6}{169}\right)}\]Taking the root of 169,\[\Large\bf\sf 13\sqrt{1+\frac{6}{169}}\]I dunno I'm just trying to think of an idea for estimating. This doesn't exactly get us to a decimal value.. hmm

OpenStudy (anonymous):

idk either..lol

OpenStudy (anonymous):

ankit042?